OpenStudy (dbzfan836):

I need help with a 2 step equation. [x+4]/2=7

OpenStudy (dbzfan836):

is the first step [x+4]/2*2=7*2, and the second x+4-4=14-4, so x=10?

OpenStudy (mokeira):

exactly...you are correct

OpenStudy (dbzfan836):

Wow, that was fast XD

OpenStudy (dbzfan836):

alright, i have another to ask. yet i have no clue how to solve

OpenStudy (dbzfan836):

1/3[2x-8]=4

OpenStudy (mokeira):

do you want to solve it or get two step explanation on how to solve?

OpenStudy (dbzfan836):

I want an explanation on how, yet i would like to try to calculate it on my own

OpenStudy (dbzfan836):

I just dont know how to solve [the order of it and some other things]

OpenStudy (mokeira):

ok Step 1 multiply both sides by 3 Step2 add 8 to both sides Step 3 divide both sides by 2

OpenStudy (dbzfan836):

oh, so i made a mistake. I thought that you multiply both sides by 1/3

OpenStudy (mokeira):

no...because you want to remove the denominator

OpenStudy (anonymous):

x + 4 = 7 x 2 x + 4 = 14 x = 14 - 4 x = 10

OpenStudy (dbzfan836):

alright step 1: 1/3*3[2x-8]=4*3 2x-8=12

OpenStudy (dbzfan836):

step 2: 2x-8+8=12+8 2x=20

OpenStudy (mokeira):

thats right

OpenStudy (dbzfan836):

step 3: 2x/2=20/2 x=10

OpenStudy (dbzfan836):

well thats all, folks -porky pig